Which expression has the same value as 25^1/2?
A. 125^2/3
B. -5
C. 25^ 1/3
D. sqrt(-5)^2

Answer by addingup(3677)

25^1/2 = 5 which is the same result you get with D, sqrt(-5)^2
NOTE: remember any number squared is positive, in the case of D:
(-5)^2 = -5 x -5 = +25 and the square root of 25 is 5